5.1. Suppose that the price of whiskey increases from R100 to R150 a bottle and as a result the quantity demanded decreases from 1 100 bottles to 800 bottles.

5.1.1. Use the ARC (midpoint) formula to calculate the price elasticity of demand for whiskey. (3)

5.1.2. Based on the calculated elasticity value in 5.1.1, indicate whether the demand in the market for whiskey is elastic or inelastic. (1)

5.1.3. Based on your answer in 5.1.2, illustrate the elasticity of demand in the market for whiskey. Clearly indicate the correct percentage changes in price and quantity on the elasticity graph. (4)

5.1.4. Explain how producers could increase total revenue given the calculated elasticity coefficient.


1
Expert's answer
2021-04-07T10:24:20-0400

Answer

5.1.1

"Arc elasticity of demand"

"Midpoint Q_{d}= \frac{800+1100} {2} = 950"


"Midpoint price=\frac{150+100}{2}=125"


"Percentage change of quantity demanded= \frac{1100-800}{950}=0.32"


"Percentage change in price= \frac{100-150}{125}=0.4"


"Arc percentage in demand= \frac{0.32}{-0.4}=0.8"


5.1.2

the demand in the market is inelastic since the formula is creating a value less than 1. Again, it the inelastic demand is as a result of the small change in quantity demanded due to change in price.

5.1.3




5.1.4

How producer increase total revenue given the calculated elasticity coefficient

The producers could increase revenue by lowering the price from R100 downwards hence the total revenue will increase as price increase.